OpenStudy (anonymous):

what is the square roots of 50 and 8

OpenStudy (leader):

\[\sqrt{50}=7.07106781187\]

OpenStudy (leader):

\[\sqrt{8}=2.82842712475\]

OpenStudy (leader):

Does that help you @ziam

OpenStudy (jhannybean):

\[\ \sqrt{50} = \sqrt{25 \cdot 2} = \sqrt{\color{red}{5\cdot 5}\cdot 2}=5\sqrt{2} \]\[\ \sqrt{8} = \sqrt{\color{red}{2 \cdot 2}\cdot 2} = 2\sqrt{2}\]

OpenStudy (anonymous):

the answer in square root